OpenStudy (anonymous):

@epoweritheta The square of a number decreased by 25 is -9. A. Find the x-intercepts. B. Find the vertex of the graph. C. What shape is the graph? Describe what you see.

OpenStudy (anonymous):

I need help with part B and C

OpenStudy (epoweritheta):

x^2-25=-9 x^2=16 ; x=+4;x=-4 what graph do u want me to draw ??

OpenStudy (anonymous):

Idk thy say find the vertex of the graph, Im not sure what's supposed to be graphed

OpenStudy (anonymous):

@phi

OpenStudy (phi):

the vertex is the "bottom" of the graph. if you graph y= x^2-16 you get a u shaped curve (a parabola)

OpenStudy (phi):

This question is very vague strictly speaking, "The square of a number decreased by 25 is -9." is not going to give us a curve we have to "read their mind" to guess they want us to graph y = x^2 -16 (and that is just a guess as to what they want)

OpenStudy (anonymous):

Is the vertex an exact point on the graph, or are they asking for the definition?

OpenStudy (phi):

they want the (x,y) point of the vertex here is a graph of y= x^2 - 16

OpenStudy (anonymous):

So that would be (0, -16)

OpenStudy (phi):

yes.

OpenStudy (anonymous):

Thank you

OpenStudy (phi):

I assume you can do part C. It's a parabola